OpenStudy (atlchic):

Solve, r-1/r+1=3/5. Show work.

OpenStudy (anonymous):

(r-4)/(r+1) = 0 (r-4)=0 r=4

OpenStudy (precal):

 cross multiply

OpenStudy (precal):